OpenStudy (anonymous):

For the linear equation, x + y - 1 = 0, determine a corresponding value of y for the x-value of -6. Type a numerical answer in the space provided. Do not type spaces in your answer.

OpenStudy (jim766):

x + y - 1 = 0 they are giving you the x, so you can find the y. just substitute it in -6 + y - 1 = 0 now can you get y alone?

OpenStudy (jim766):

first combine like terms -7 + y = 0 now do you know how to get y alone?